The parabola given by the equation y=−x^2+12x−34 has its vertex at (h,k) for:
h=
and
k=

Answer by stanbon(75887)

The parabola given by the equation y=−x^2+12x−34 has its vertex at (h,k) for:
-----
Complete the square:
-x^2 + 12x = y+34
-----
-(x^2-12x+36) = y + 34 - 36
----
-(x-6)^2 = y-2
-----
h= 6
and
k= 2
============
Cheers,
Stan H.
============